# Neural Deconstruction Search for Vehicle Routing Problems

### v1

@@ Appendix A Training Curves
Figure 4 presents the training curves for all experiments conducted across the three problem types and four problem sizes. Note that the training of the models for $N$=$2000$ is warm-started using the model weights from $N$=$1000$ after 1,500 epochs.
$N$=$100$
$N$=$500$
$N$=$1000$
$N$=$2000$
Figure 4: Training curves.
Appendix B Visualizations of Policy Rollouts
@@
Figures 5, 6, and 7 show visualizations of different policy rollouts for the CVRP, PCVRP, and VRPTW, respectively. For each problem, we display two different instances, and for each instance, six rollouts are shown. Customers selected for deconstruction are circled in red. We note that the nodes selected for each deconstruction differs, sometimes significantly, allowing NDS to try out a variety of options in each iteration.
@@
Figure 5: Rollouts for two selected instances for the CVRP with $N$=$100$ (best viewed in color).
@@
Figure 6: Rollouts for two selected instances for the PCVRP with $N$=$100$ (best viewed in color).
@@
Figure 7: Rollouts for two selected instances for the VRPTW with $N$=$100$ (best viewed in color).

### v2

@@ Appendix B Training Curves
Figure 5 presents the validation performance throughout training for all experiments conducted across the three problem types and four problem sizes. Note that the training of the models for NN=$2000$ is warm-started using the model weights from NN=$1000$ after 1,500 epochs.
Figure 5: Validation performance throughout the training process.
Appendix C Visualizations of Policy Rollouts
@@
Figures 6, 8, and 7 show visualizations of different policy rollouts for the CVRP, PCVRP, and VRPTW, respectively. For each problem, we display two different instances, and for each instance, six rollouts are shown. Customers selected for deconstruction are circled in red. We note that the nodes selected for each deconstruction differs, sometimes significantly, allowing NDS to try out a variety of options in each iteration.
@@
Figure 6: Rollouts for two selected instances for the CVRP with NN=$100$ (best viewed in color).
@@
Figure 7: Rollouts for two selected instances for the VRPTW with NN=$100$ (best viewed in color).
@@
Figure 8: Rollouts for two selected instances for the PCVRP with NN=$100$ (best viewed in color).
Appendix D Performance Across Different Runtime Limits
We evaluate the performance of NDS and SISRs under varying runtime limits. The results are based on the same trained models and evaluation setup described in Section 4.4, but with test sets reduced to 100 instances each to manage computational costs. The hyperparameters remain as described in Section 4.4, except for the number of augmentations for the PCVRP. Specifically, for the PCVRP with 1000 customers, 256 augmentations are used for runtime limits of 240 seconds or more. For the PCVRP with 2000 customers, 64 augmentations are applied for runtime limits of 240 seconds or less, and 256 augmentations for a runtime limit of 16 minutes. The runtime limits in this experiment are selected based on instance size, ranging from 2 seconds for instances with 100 customers to 16 minutes for instances with 2000 customers.
Figure 9 presents the results across all problems and problem sizes. Overall, NDS significantly outperforms SISRs on all problems. SISRs only surpasses NDS in a few cases under very short runtime limits. The performance gap between NDS and SISRs appears to depend more on problem size than on problem type. For example, on instances with 100 customers, SISRs struggles to achieve the solution quality that NDS attains in just 2 seconds, even after 60 seconds of runtime.
Figure 9: Performance of NDS and SISRs across different runtime limits.
Appendix E Impact of Hyperparameters on Search Performance
We evaluate the impact of key hyperparameters in the ASA search procedure. To manage computational costs, this experiment is conducted solely on the CVRP with 500 nodes. We perform multiple evaluation runs on the test set, varying the number of augmentations $A$, the number of rollouts $K$, and the threshold factor $\delta$, while keeping all other parameters at their default values as specified in Section 5.2.
The results, presented in Figure 11, indicate that values of $A$ equal or greater than 8 yield the best performance. The number of rollouts $K$ has little impact on performance within the tested range of 100 to 300. For the threshold factor $\delta$, values above 15 yield consistently strong results. Overall, the search procedure demonstrates robustness to hyperparameter choices, suggesting that fine-tuning is not critical for achieving good performance.
Figure 10: Search performance on the CVRP500 for different hyperparameter setting.
Appendix F Ablation Study: Initial Improvement Steps
We assess the impact of initial solution improvement through an ablation study. To this end, we train models on problem instances with 500 nodes, omitting any solution improvement steps before the main training phase (lines 10–17 in Algorithm 1). As a result, the model also learns to refine low-quality start solutions.
Figure 11 compares validation performance during training with and without the initial improvement steps. The results reveal substantial performance differences, with training runs that include initial improvement significantly outperforming those without, highlighting the effectiveness of this technique. We hypothesize that initial improvement is beneficial because learning directly from poor-quality start solutions provides limited value; improving them is trivial, whereas most of the search process focuses on refining higher-quality solutions.
Figure 11: Validation performance for training runs with and without initial solution improvement steps.
Appendix G Generalization to Smaller and Larger Instances
We also evaluate the generalization ability of NDS to larger and smaller instances by using models trained on instances with 500 customers to solve instances with 100, 1000, and 2000 customers. In this experiment, all hyperparameters are set identically to those in the main experiment.
The results, presented in Table 4, also include in-distribution performance from the main experiment, where instance size-specific models are used, to allow for an easier comparison. The findings indicate that models trained on 500 customer instances perform well on instances with 100 and 1000 customers, with no significant degradation in performance. However, on instances with 2000 customers, a slight drop in performance is observed, particularly for the CVRP.
Table 4: Out-of-distribution (OOD) performance of models trained on instances with 500 nodes, evaluated on both larger and smaller instances, compared to the in-distribution (ID) performance of size-specific models.
N=100 N=1000 N=2000
(Smaller) (Larger) (Much Larger)
Obj. Gap Time Obj. Gap Time Obj. Gap Time
CVRP HGS 15.57 - 5 41.51 - 121 57.38 - 241
NDS (ID) 15.57 0.04% 5 41.11 -0.90% 120 56.00 -2.34% 240
NDS (OOD) 15.59 0.13% 5 41.32 -0.45% 120 57.63 0.43% 240
VRPTW PyVRP-HGS 12.98 - 5 90.35 - 120 173.46 - 240
NDS (ID) 12.95 -0.19% 5 87.54 -3.14% 120 167.48 -3.50% 240
NDS (OOD) 12.97 -0.00% 5 87.54 -3.11% 120 167.76 -3.28% 240
PCVRP PyVRP-HGS 10.11 - 5 84.91 - 120 165.56 - 240
NDS (ID) 9.90 -2.07% 5 80.99 -4.71% 121 158.09 -4.60% 241
NDS (OOD) 9.96 -1.50% 5 81.05 -4.55% 120 160.20 -3.24% 240